OpenStudy (anonymous):

Select the equation of the line in slope-intercept form: 4x – 9y = –36 Can someone help me solve this? :)

OpenStudy (anonymous):

ok slope intercept form is y=mx+b have you seen this before?

OpenStudy (anonymous):

4x – 9y = –36 9y=4x + 36 y=4x/9 + 36/9 y=4x/9 + 4

OpenStudy (anonymous):

Yes. I just don't know how to plug MY equation in with that equation. T_T

OpenStudy (anonymous):

you want to isolate your y variable. to do that you need to subtract 4x from both sides of the equation. you would then have: -9y=-4x-36 next you need to divide both sides by -9 and this would be your final answer in slope intercept form y=mx+b \[y=\frac{ 4 }{ 9 }x+4\]